### v2

# CHASE: 3D-Consistent Human Avatars with Sparse Inputs via Gaussian Splatting and Contrastive Learning

Abstract
Recent advancements in human avatar synthesis have utilized radiance fields to reconstruct photo-realistic animatable human avatars. However, both NeRFs-based and 3DGS-based methods struggle with maintaining 3D consistency and exhibit suboptimal detail reconstruction, especially with sparse inputs. To address this challenge, we propose CHASE, which introduces supervision from intrinsic 3D consistency across poses and 3D geometry contrastive learning, achieving performance comparable with sparse inputs to that with full inputs. Following previous work, we first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to coordinate the movements of individual Gaussians during animation, reconstructing basic avatar with coarse 3D consistency. To improve 3D consistency under sparse inputs, we design Dynamic Avatar Adjustment(DAA) to adjust deformed Gaussians based on a selected similar pose/image from the dataset. Minimizing the difference between the image rendered by adjusted Gaussians and the image with the similar pose serves as an additional form of supervision for avatar. Furthermore, we propose a 3D geometry contrastive learning strategy to maintain the 3D global consistency of generated avatars. Though CHASE is designed for sparse inputs, it surprisingly outperforms current SOTA methods in both full and sparse settings on the ZJU-MoCap and H36M datasets, demonstrating that our CHASE successfully maintains avatar’s 3D consistency, hence improving rendering quality. The code will be made available.
footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
Introduction
Photorealistic rendering and animation of human bodies is a crucial research area with wide-ranging applications in AR/VR, visual effects, visual try-on, and movie production (Zackariasson and Wilson 2012; Healey, Wang, and et al 2021). Early works (Mildenhall et al. 2021; Oechsle, Peng, and Geiger 2021; Niemeyer et al. 2020; Peng et al. 2021b; Wang et al. 2022a; Sitzmann et al. 2021) for creating human avatars relied on capturing high-quality data through multi-camera setups, which is extensive computing and needs lots of manual effort. While these methods excel at creating a single scene/object from sufficient input views, it is very challenging for them to generalize to new scenes/objects with few samples (Kwon et al. 2024).
Recent advancements have explored using neural radiance fields (NeRF) for modeling 3D human avatars (Mildenhall et al. 2021; Wang et al. 2022a), typically employing parametric body models to model deformations. Some methods (Chen et al. 2022; Zhao et al. 2022) use human template models to facilitate generalizable and robust synthesis. Though NeRF-based methods have made significant progress in generalizable human rendering, they are less efficient to train and render due to their computationally intensive per-pixel volume rendering process.
Point-based rendering (Zheng et al. 2023) has emerged as an efficient alternative to NeRFs, offering significantly faster rendering. The recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) gains popularity for its efficient rasterization-based rendering speed. Numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatars (Lei et al. 2024; Moreau et al. 2024; Shao et al. 2024; Hu et al. 2024b; Kocabas et al. 2024; Wang et al. 2024a; Qian et al. 2024; Hu et al. 2024a). However, these methods often face challenges in maintaining 3D consistency and producing high-quality reconstructions, particularly with sparse inputs.
To address the aforementioned issues, we propose CHASE, which is capable of reconstructing 3D Consistent Human Avatars with Sparse inputs via Gaussian Splatting and contrastivE learning. We first integrates a skeleton-driven rigid deformation, and a non-rigid cloth dynamics deformation to create human avatar with coarse 3D consistency. To enhance 3D consistency under sparse inputs, we utilize the intrinsic 3D consistency of images across different poses within the same person. Specifically, for each training pose/image, we select a similar pose/image from the dataset and then adjust the deformed Gaussians using the proposed Dynamic Avatar Adjustment (DAA), an explicit point-based control graph adjustment strategy, to the selected similar pose. Then we minimize differences between the rendered image of the adjusted Gaussians and the image corresponding to the selected similar pose, which serves as an additional form of supervision for human avatar. Additionally, we employ 3D geometry contrastive learning, utilizing features from a 3D feature extractor, to further enhance the global 3D consistency of generated human avatars. We conduct extensive experiments on the ZJU-MoCap data (Peng et al. 2021c), H36M (Ionescu et al. 2013) and find that CHASE not only achieves better performance in sparse inputs setting but also outperforms other SOTAs in full setting.
In summary, our work makes the following contributions:
  • We propose an explicit point-based control graph adjustment strategy, which introduces a novel 2D image supervision to 3D human body modeling, enhancing the 3D consistency of human avatars.
  • We propose 3D geometry contrastive learning to enforce consistency across different representations of the same pose and facilitate effective 3D global understanding.
  • Extensive experiments show that our CHASE achieves SOTA performance quantitatively and qualitatively under full and sparse settings.
Related Work
Contrastive Representation Learning
Contrastive Representation Learning is one of the mainstream self-supervised learning paradigms (Hadsell, Chopra, and LeCun 2006), which learns potential semantics from constructed invariance or equivariance (Rumen et al. 2022). In 3D, PointContrast (Xie et al. 2020) proposes geometric augmentation to generate positive and negative pairs. CrossPoint (Afham et al. 2022) uses both inter- and intra-modal contrastive learning. PointCLIP (Zhang et al. 2022) achieves image-point alignment by projecting point clouds onto 2D depth images. RECON (Qi et al. 2023) focuses on single- and cross-modal contrastive learning through discriminative contrast (Khosla et al. 2020) or global feature alignment (Radford et al. 2021). Our CHASE introduces a 3D geometry contrastive learning method across different poses to enforce consistency across different representations of the same pose.
3D Editing and Deformation
Traditional deformation methods in computer graphics are typically based on Laplacian coordinates (Gao et al. 2019; Sorkine and Alexa 2007; Sorkine 2005), Poisson equations (Yu et al. 2004), and cage-based methods (Yifan et al. 2020). However, these methods often rely on implicit and computationally expensive NeRF-based approaches.
Numerous works (Chen et al. 2024; Wang et al. 2024b) have proposed techniques for editing 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023). SuGaR(Guédon and Lepetit 2024) introduces a mesh extraction method that produces meshes from 3DGS, which can then be edited. SC-GS (Huang et al. 2024) proposes deforming Gaussians by transferring the movement of control points. Our CHASE employs a novel explicit point-based control graph deformation strategy, which is more intuitive and efficient.
3D Human Modeling
Since the high-quality rendering achieved by the seminal work Neural Radiance Fields (NeRF)(Mildenhall et al. 2021), there has been a surge of research on neural rendering for human avatars(Liu et al. 2021; Li et al. 2023; Peng et al. 2021c). Although NeRF is designed for static objects, HumanNeRF (Weng et al. 2022) extends NeRF to enable capturing dynamic human motion using just a single monocular video. Neural Body (Peng et al. 2021c) associates a latent code with each SMPL (Loper et al. 2015) vertex to encode appearance, which is then transformed into observation space based on the human pose. Furthermore, Neural Actor (Liu et al. 2021) learns a deformable radiance field with SMPL (Loper et al. 2015) as guidance and utilizes a texture map to improve the final rendering quality. Posevocab (Li et al. 2023) designs joint-structured pose embeddings to encode dynamic appearances under different key poses, allowing for more effective learning of joint-related appearances. However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
Point-based rendering (Zheng et al. 2023) has proven to be an efficient alternative to NeRFs for fast inference and training. Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS)(Kerbl et al. 2023) models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending. Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatar reconstruction(Lei et al. 2024; Qian et al. 2024; Hu et al. 2024b; Kocabas et al. 2024). Human Gaussian Splatting (Moreau et al. 2024) showcases 3DGS as an efficient alternative to NeRF. SplattingAvatar (Shao et al. 2024) and GomAvatar (Wen et al. 2024) extend lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh. However, these methods struggle to maintain 3D consistency and produce low-quality reconstructions when applied to human avatar creation with only sparse inputs. Our CHASE introduces a novel 2D image supervision to 3D human body modeling and 3D geometry contrastive learning, enhancing the 3D consistency of human avatars.
Figure 1: CHASE Framework. We first initialize 3D Gaussians in canonical space by randomly sampling 50k points on the SMPL mesh surface. Then, we integrate a rigid human articulation and a non-rigid deformation neural field to deform the 3D Gaussians in canonical space ${\mathcal{G}_{c}}$ to the observation space ${\mathcal{G}_{o}}$. Next, we select similar poses/images from the dataset for each training pose/image and then adjust the deformed Gaussians ${\mathcal{G}_{o}}$ to the similar pose ${\mathcal{G}_{a}}$ using Dynamic Avatar Adjustment (DAA). Minimizing the differences between the rendered adjusted Gaussians ${\mathcal{G}_{a}}$ and the selected similar images $x_{i}^{a}$ serves as an additional supervision. Furthermore, we propose a 3D geometry contrastive learning, which involves comparing features from a 3D feature extractor to improve the avatar’s global 3D consistency. Negative pairs consist of the features of the deformed Gaussians ${\mathcal{G}_{o}}$ and the adjusted Gaussians ${\mathcal{G}_{a}}$. In contrast, positive pairs include the features of ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the pose adjustments seen in ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{a}}$.
Preliminaries
SMPL (Loper et al. 2015). The SMPL model is a pre-trained parametric human model representing body shape and pose. In SMPL, body shape and pose are controlled by pose and shape. In this work, we apply the Linear Blend Skinning (LBS) algorithm used in SMPL to transform points from a canonical space to a posed space.
LBS (Sumner, Schmid, and Pauly 2007). Linear Blend Skinning (LBS) is a weight-based technique that associates each vertex with one or more joints and uses weight values to describe the influence of each joint on the vertex. Vertex deformation is calculated by linearly interpolating transformations on the associated joints: $\mathcal{X}_{v}^{{}^{\prime}}=\sum_{j=1}^{J}w_{j}(\mathcal{X}_{v})B_{j}\mathcal{X}_{v}$, where $J$ represents the number of joints, $N$ represents the number of vertices, $\mathcal{X}_{v}^{{}^{\prime}}\in\mathbb{R}^{N\times 3}$ is the new position of the skinned vertex, $w\in\mathbb{R}^{N\times J}$ is the skinning weight matrix, $B\in\mathbb{R}^{J\times 4\times 4}$ is the affine transformation matrix of each joint representing rotation and translation, (i.e. bone transforms) and $\mathcal{X}_{v}\in\mathbb{R}^{N\times 3}$ is the original mesh vertex position.
3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean. The value at point $\mathcal{X}$ is:
$G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ (1)
For differentiable optimization, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathcal{S}$ and a rotation matrix $\mathcal{R}$, such that $\Sigma=\mathcal{R}\mathcal{S}\mathcal{S}^{T}\mathcal{R}^{T}$. $\mathcal{S}$ and $\mathcal{R}$ are stored as the diagonal vector $s\in\mathbb{R}^{N\times 3}$ and a quaternion vector $r\in\mathbb{R}^{N\times 4}$, respectively.
In rendering novel views, differential splatting as introduced by (Yifan et al. 2019), involves using a viewing transform $W$ and the Jacobian matrix $J$ of the affine approximation of the projective transformation to compute the transformed covariance matrix: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color and opacity at each pixel are computed from the Gaussian’s representations: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ The blending of $N$ ordered points overlapping a pixel is given by the formula:
$\mathcal{C}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ (2)
where $c_{i}$, $\alpha_{i}$ represent the density and color of this point computed by a 3D Gaussian $G$ with covariance $\Sigma$ multiplied by an optimizable per-point opacity and SH color coefficients.
Method
We illustrate the pipeline of our CHASE in Fig. 1. The inputs include images $X=\{x_{i}\}^{N}_{i=1}$ obtained from monocular videos, fitted SMPL parameters $P=\{p_{i}\}^{N}_{i=1}$, and foreground masks $M=\{m_{i}\}^{N}_{i=1}$ of images. CHASE optimizes 3D Gaussians in canonical space, which are then be deformed to match the observation space and be rendered with a given camera view.
Non-rigid and Rigid Deformation
Inspired by (Weng et al. 2022; Qian et al. 2024), we deform 3D Gaussians from canonical space ${\mathcal{G}_{c}}$ to observation space ${\mathcal{G}_{o}}$ by integrating a rigid articulation with a non-rigid transformation. We employ a non-rigid deformation network that takes the canonical positions $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}c}$ and a pose latent code which encodes SMPL pose $p_{i}$ using a lightweight hierarchical pose encoder (Mihajlovic et al. 2021) as input. The network then outputs the offsets for various parameters of the 3D Gaussians $\mathcal{G}_{c}$: $\Delta{(\mathcal{X},\mathcal{C},\alpha,s,r)}$. The canonical Gaussians are deformed by:
$\mathcal{X}_{d}$ $=\mathcal{X}_{c}+\Delta\mathcal{X},\mathcal{C}_{d}=\mathcal{C}_{c}+\Delta\mathcal{C},$ (3)
$\alpha_{d}$ $=\alpha_{c}+\Delta\alpha,s_{d}=s_{c}\cdot\exp(\Delta s),$ (4)
$r_{d}$ $=r_{c}\cdot[1,\Delta r_{1},\Delta r_{2},\Delta r_{3}],$ (5)
where quaternion multiplication $\cdot$ corresponds to multiplying the rotation matrices. With $[1,0,0,0]$ as the identity rotation, $r_{d}=r_{c}$ when $\delta{r}=\mathbf{0}$, thus keeping the original orientation.
We further apply a LBS-based rigid transformation to map the non-rigidly deformed 3D Gaussians ${\mathcal{G}_{d}}$ to the observation space ${\mathcal{G}_{o}}$. This transformation utilizes LBS weights predicted by a Skinning MLP $f_{\theta_{r}}$. This process aligns the Gaussians with the target pose in ${\mathcal{G}_{o}}$:
$T=\sum_{j=1}^{J}f_{\theta_{r}}(\mathcal{X}_{d})_{j}B_{j},\mathcal{X}_{o}=T\mathcal{X}_{d},$ (6)
$\mathcal{R}_{o}=T_{1:3,1:3}\mathcal{R}_{d},$ (7)
where $\mathcal{R}$ is the matrix representations of rotation.
Figure 2: For each training pose/image, we select similar poses/images from the dataset and then adjust the deformed Gaussians using Dynamic Avatar Adjustment (DAA). By minimizing the discrepancy between the rendered image of the adjusted avatar and the selected similar pose image, we introduce an additional supervision, thereby enhancing the creation of animatable avatars.
Dynamic Avatar Adjustment
To address extremely sparse inputs, we leverage the intrinsic 3D consistency of human avatars across different poses/images, as shown in Fig. 2. Specifically, for each training pose/image, we select a similar pose $p^{a}_{i}$ with its paired image $x_{i}^{a}$ from the dataset and then use a dense motion field $F_{adj}$ as an additional adjustment to transform deformed Gaussians ${\mathcal{G}_{o}}$ into adjusted Gaussians ${\mathcal{G}_{a}}$, aligning them with the selected pose/image ($p_{i}^{a}$/$x_{i}^{a}$). In this way, we successfully introduce an additional 2D image supervision, improving the 3D consistency of human avatars.
To achieve precise control of the 3D Gaussians, we sample 6,890 points from the SMPL model (Loper et al. 2015) as our sparse control points in canonical space. Then, we obtain the dense motion field using LBS by locally inheriting the LBS weights from neighboring control points. Specifically, for each 3D Gaussian, we use the k-nearest neighbor (KNN) search to find its nearest neighboring control points in canonical space. The entire adjustment process is as:
$w=w_{smpl}[\text{KNN}(xyz_{cano},xyz_{smpl})],$ (8)
$T_{o}=\sum_{j=1}^{J}w_{j}B_{oj},T_{o}^{{}^{\prime}}=\sum_{j=1}^{J}w_{j}B_{oj}^{{}^{\prime}}.$ (9)
Here, $w_{smpl}$ denotes the LBS weights of the sparse control points, and $T_{o}(B_{o})$ and $T_{o}^{\prime}(B_{o}^{\prime})$ represent the rigid transformations (bone transformations) from canonical space $\mathcal{G}_{c}$ to deformed Gaussians $\mathcal{G}_{o}$, and to the Gaussians with the selected similar pose $\mathcal{G}_{o}^{\prime}$, respectively. We then obtain $F_{adj}$, which transforms the deformed Gaussians $\mathcal{G}_{o}$ into adjusted Gaussians $\mathcal{G}_{a}$, aligning them with the selected pose $p_{i}^{a}$ as:
$F_{adj}$ $=F_{o^{\prime}}F_{o}^{-1}.$ (10)
We adjust the deformed Gaussians ${\mathcal{G}_{o}}$ to adjusted Gaussians ${\mathcal{G}_{a}}$ by by adjusting its position and rotation as follow:
$\mathcal{X}_{a}$ $=F_{adj}\mathcal{X}_{o},$ (11)
$\mathcal{R}_{a}$ $=F_{adj1:3,1:3}\mathcal{R}_{o}.$ (12)
3D Geometry Contrastive learning
Inspired by the success of contrastive learning in image processing and static point cloud analysis, we advocate for adopting 3D geometry contrastive learning to ensure 3D consistency during animation. To this end, we treat the 3D Gaussians as a 3D point cloud and use DGCNN (Wang et al. 2019) as the feature extractor. The point cloud feature extractor processes the positions of the 3D Gaussians in the observation space ${\mathcal{G}_{o}}$, the adjusted Gaussians ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the selected pose $p_{i}^{a}$, and outputs their features, creating intermediate graph features to capture global geometric information better. The feature vectors are projected into an invariant space. We denote the projected features of ${\mathcal{G}_{o}}$, ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$ as $f_{o}$, $f_{a}$, and $f_{o}^{\prime}$, respectively.
In the invariant space, we aim to maximize the similarity between $f_{a}$ and $f_{o}^{\prime}$, denoted as $\text{D}_{positive}$, and minimize the similarity between $f_{a}$ and $f_{o}$, denoted as $\text{D}_{negative}$. Therefore, we compute the 3D geometry contrastive loss $\mathcal{L}_{contrastive}$ as:
$\text{D}_{positive}$ $=\|f_{a}-f_{o}^{\prime}\|_{2},$ (13)
$\text{D}_{negative}$ $=\|f_{a}-f_{o}\|_{2},$ (14)
$\mathcal{L}_{contrastive}$ $=\max(0,\text{D}_{positive}-\text{D}_{negative}).$ (15)
Optimization
Color MLP.
Following (Qian et al. 2024), we use the inverse rigid transformation to canonicalize the viewing direction: $\hat{d}=T_{1:3,1:3}^{-1}d$, where $T$ and $d$ is the forward transformation matrix defined in LBS and viewing direction, respectively. Theoretically, canonicalizing viewing direction also promotes consistency of the specular component of canonical 3D Gaussians under rigid transformations.
Pose correction.
Following (Qian et al. 2024), SMPL (Loper et al. 2015) parameter fittings from images can be inaccurate. We additionally optimize the per-sequence shape parameter and per-frame translation, global rotation, and local joint rotations.
Loss function.
Our full loss function consists of several components: an RGB loss $\mathcal{L}_{rgb}$, a mask loss $\mathcal{L}_{mask}$, and a perceptual similarity (LPIPS) loss $\mathcal{L}_{LPIPS}$. We compute these losses on both images rendered from the deformed Gaussians $\mathcal{G}_{o}$ and the adjusted Gaussians $\mathcal{G}_{a}$ with their corresponding ground truth images. Additionally, we include a skinning weight regularization loss $\mathcal{L}_{skin}$, as well as isometric regularization losses for both position and covariance, $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$, respectively, following (Qian et al. 2024). We also incorporate a 3D geometry contrastive loss $\mathcal{L}_{contrastive}$:
$\mathcal{L}=$ $\mathcal{L}_{rgb}+\lambda_{1}\mathcal{L}_{mask}+\lambda_{2}\mathcal{L}_{LPIPS}+\lambda_{3}\mathcal{L}_{skin}+$
$\lambda_{4}\mathcal{L}_{isopos}+\lambda_{5}\mathcal{L}_{isocov}+\lambda_{6}\mathcal{L}_{contrastive},$ (16)
where $\lambda$’s are loss weights. For further details of the loss definition and respective weights, please refer to the Supp.Mat.
Figure 3: Qualitative Comparison on ZJU-MoCap (Peng et al. 2021c). We present results for full and sparse inputs (5% of the full inputs) on the ZJU-MoCap dataset. Results show that our CHASE can produce realistic details with both full and sparse inputs, while other approaches struggle to generate smooth details.
Table 1: Quantitative Results on ZJU-MoCap (Peng et al. 2021c). CHASE achieves state-of-the-art performance across every method. The best and the second best results are denoted by pink and yellow. Frames per second (FPS) is measured on an RTX 3090. We train our model on the dataset that includes only 5% of the origin data for fair quantitative comparison. The metrics are reported in the last four rows of the table. LPIPS = LPIPS $\times$ 1000.
[…]
NeuralBody (Peng et al. 2021c) 29.07 0.962 52.29 1.5
Ani-NeRF (Peng et al. 2021a) 29.17 0.961 51.98 1.1
HumanNeRF (Weng et al. 2022) 30.24 0.968 31.73 0.3
MonoHuman (Yu et al. 2023) 29.38 0.964 37.51 0.1
3DGS-Avatar (Qian et al. 2024) 30.62 0.965 30.28 50
GauHuman (Hu et al. 2024b) 30.79 0.960 32.73 180
GoMAvatar (Wen et al. 2024) 30.37 0.969 32.53 43
Ours 30.81 0.970 27.48 50
3DGS-Avatar* (Qian et al. 2024) 29.98 0.957 40.01 50
GauHuman* (Hu et al. 2024b) 30.35 0.957 35.68 180
GoMAvatar* (Wen et al. 2024) 30.01 0.958 42.88 43
Ours* 30.48 0.969 29.94 50
Table 2: Quantitative Results on H36M (Ionescu et al. 2013). Our CHASE outperforms current SOTA methods in both full and sparse settings
Training Poses Novel Poses
[…]
NARF (Noguchi et al. 2021) 23.00 0.898 22.27 0.881
NeuralBody (Peng et al. 2021c) 22.89 0.896 23.09 0.891
Ani-NeRF (Peng et al. 2021a) 23.00 0.890 22.55 0.880
ARAH (Wang et al. 2022b) 24.79 0.918 23.42 0.896
3DGS-Avatar (Qian et al. 2024) 32.89 0.982 32.50 0.983
Ours 33.29 0.984 32.93 0.982
3DGS-Avatar* (Qian et al. 2024) 32.48 0.976 32.17 0.981
Ours* 32.91 0.983 32.64 0.982
Experiment
Dataset
ZJU-MoCap (Peng et al. 2021c). This dataset features multi-view videos captured by 21 cameras, with human poses recorded using a marker-less motion capture system. For our experiments, we selected six sequences (377, 386, 387, 392, 393, 394). Following the protocol established by HumanNeRF (Weng et al. 2022) and 3DGS-Avatar (Qian et al. 2024), we use a single camera for training and the remaining cameras for evaluation. The foreground masks, camera, and SMPL parameters provided by the data set are used for evaluation purposes. We use 5% of the dataset to simulate sparse inputs.
H36M (Ionescu et al. 2013). H36M is another widely used dataset for human avatar research, comprising multi-view videos from four cameras and human poses captured via a marker-based motion capture system. It features multiple subjects performing a variety of complex actions. We conducted experiments on sequences from subjects S1, S5, S6, S7, S8, S9, and S11, selecting representative actions and dividing the videos into training and test frames. Adhering to the protocol set by ARAH (Wang et al. 2022b), we use three cameras, [54138969, 55011271, 58860488],for training and the remaining camera, [60457274],for testing, and follow their preprocessing steps. We use the SMPL parameters and foreground humans following (Peng et al. 2021a).
Comparison with State-of-the-art Methods
We compare our CHASE with various SOTA methods for human avatars, including NerF-based methods such as NeuralBody (Peng et al. 2021c), Ani-NeRF (Peng et al. 2021a), HumanNeRF (Weng et al. 2022), and MonoHuman (Yu et al. 2023), and 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024), GauHuman (Hu et al. 2024b), and GoMAvatar (Wen et al. 2024) under monocular setup on ZJU-MoCap (Peng et al. 2021c). The quantitative results are shown in Tab. 1. Overall, our proposed CHASE achieves the best performance in terms of PSNR, SSIM, and LPIPS with both full and sparse inputs. Our CHASE shows only a small performance drop when using only 5% of the data and significantly outperforms SOTA methods on LPIPS. It is evidence that our method successfully maintains 3D consistency even with sparse inputs.
Qualitative comparisons on novel view synthesis are shown in Fig. 3. We observe that our method preserves more details compared to other SOTA methods. They often struggle to maintain 3D consistency and deliver suboptimal detail reconstruction in human avatar modeling, particularly when only sparse inputs are available.
For H36M (Ionescu et al. 2013), we report the quantitative results against NerF-based methods such as NARF (Noguchi et al. 2021), NeuralBody (Peng et al. 2021c), Ani-NeRF (Peng et al. 2021a) and ARAH (Wang et al. 2022b), and 3DGS-based methods such as 3DGS-Avatar (Qian et al. 2024) in Tab. 2. Our CHASE significantly outperforms these methods with sparse inputs, showing that our CHASE generalizes well to novel poses with sparse inputs and reconstructs human avatars with better appearance and geometry detail. For qualitative comparisons on novel pose synthesis, please refer to the Supp.Mat.
Table 3: Ablation Study on ZJU-MoCap (Peng et al. 2021c). We both show the result from full input (top group) and sparse input (bottom group).
[…]
w/o non-rigid 30.32 0.968 30.41 50
w/o contrastive 30.76 0.970 27.58 50
pointnet 30.75 0.970 27.74 50
w/o DAA 30.78 0.970 27.83 50
Full model 30.81 0.970 27.48 50
w/o con 30.33 0.968 29.96 50
w/o DAA 30.42 0.968 30.17 50
Full model 30.41 0.969 29.94 50
Figure 4: Ablation Study on Dynamic Avatar Adjustment, which enhances multi-view consistency, hence improving rendering quality.
Figure 5: Ablation Study on 3D geometry contrastive learning, which removes the artifacts on highly articulated poses..
Ablation Study
In this section, we conduct ablation experiments using the ZJU-MoCap (Peng et al. 2021c) dataset with both full inputs and sparse inputs to evaluate the effectiveness of our proposed modules. We also conduct experiments to evaluate different backbones in 3D geometry contrastive learning. Notably, our CHASE maintains 3D consistency effectively without increasing any extra inference time.
Non-rigid Deformation. Non-rigid Deformation is designed for complex cloth deformation. As shown in Tab. 3, non-rigid deformation is required to achieve optimal performance, demonstrating non-rigid regions are well rendered.
Dynamic Avatar Adjustment. As shown in Tab.3, incorporating Dynamic Avatar Adjustment (DAA) results in our full model outperforming the baseline in terms of LPIPS. We argue that LPIPS is particularly informative compared to other metrics because it addresses the challenge of accurately reproducing ground-truth appearance from novel views in a monocular setting(Qian et al. 2024; Yang et al. 2024). Fig. 4 illustrates that Dynamic Avatar Adjustment serves as an effective 2D image supervision technique for 3D human body modeling, enhancing 3D consistency and reducing artifacts while improving multi-view consistency.
3D geometry contrastive learning. Our core idea is that cross-3D-modal contrastive learning can facilitate communication between 3D models for obtaining powerful representations. To verify this, we further do ablation studies to show qualitative comparisons in Tab. 3 and Fig. 5. We can find the full model (w/ contrastive) preserves finer details and provides a more realistic and detailed reconstruction of clothing and other complex surfaces, demonstrating that 3D Geometry Contrastive Learning enhances 3D consistency.
Backbone for 3D contrastive learning. In Tab. 3, We show the ablation study on different backbones, including PointNet (Qi et al. 2017) (pointnet) and DGCNN (Wang et al. 2019) (full model). This indicates that DGCNN captures local geometric details better by building a dynamic graph structure, while PointNet relies on global feature learning, which may cause some local information to be missing.
Conclusion
In this paper, we present CHASE, a 3D-consistent human modeling framework utilizing Gaussian Splatting with both full and sparse inputs. We first integrates a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to create human avatar. To improve 3D consistency under sparse inputs, we use the intrinsic 3D consistency of images across poses. For each training image, we select similar pose/image from the dataset and adjust the deformed Gaussians to selected pose by Dynamic Avatar Adjustment (DAA). Minimizing the difference between the image rendered by adjusted Gaussians and image paired with selected similar pose serves as an additional supervision, hence enhancing the 3D consistency of human avatars. Furthermore, to enforce global 3D consistency across different representations of the same pose, we propose a 3D geometry contrastive learning. Extensive experiments on two popular datasets demonstrate that CHASE not only achieves superior fidelity in generating human avatars compared to current SOTA methods but also excels in handling both monocular and sparse input scenarios. We hope that our method could foster further research in high-quality clothed human avatar synthesis from monocular views.
Limitations. 1). CHASE lacks the capability to extract 3D meshes. Developing a method to extract meshes from 3D Gaussians is an important direction for future research. 2). Our CHASE need to find similar pose from dataset. When it comes to monocular videos featuring large-scale movements and diverse actions, it is hard to find pairs that qualify the requirements, which may impact performance.

### v3

@@
 Abstract
-Recent advancements in human avatar synthesis have utilized radiance fields to reconstruct photo-realistic animatable human avatars.
-However, both NeRFs-based and 3DGS-based methods struggle with maintaining 3D consistency and exhibit suboptimal detail reconstruction, especially with sparse inputs.
-To address this challenge, we propose CHASE, which introduces supervision from intrinsic 3D consistency across poses and 3D geometry contrastive learning, achieving performance comparable with sparse inputs to that with full inputs.
-Following previous work, we first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to coordinate the movements of individual Gaussians during animation, reconstructing basic avatar with coarse 3D consistency.
-To improve 3D consistency under sparse inputs, we design Dynamic Avatar Adjustment(DAA) to adjust deformed Gaussians based on a selected similar pose/image from the dataset.
-Minimizing the difference between the image rendered by adjusted Gaussians and the image with the similar pose serves as an additional form of supervision for avatar.
-Furthermore, we propose a 3D geometry contrastive learning strategy to maintain the 3D global consistency of generated avatars.
-Though CHASE is designed for sparse inputs, it surprisingly outperforms current SOTA methods in both full and sparse settings on the ZJU-MoCap and H36M datasets, demonstrating that our CHASE successfully maintains avatar’s 3D consistency, hence improving rendering quality.
-The code will be made available.
-footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
-Introduction
-Photorealistic rendering and animation of human bodies is a crucial research area with wide-ranging applications in AR/VR, visual effects, visual try-on, and movie production (Zackariasson and Wilson 2012; Healey, Wang, and et al 2021).
-Early works (Mildenhall et al.
-2021; Oechsle, Peng, and Geiger 2021; Niemeyer et al.
-2020; Peng et al.
-2021b; Wang et al.
-2022a; Sitzmann et al.
-2021) for creating human avatars relied on capturing high-quality data through multi-camera setups, which is extensive computing and needs lots of manual effort.
-While these methods excel at creating a single scene/object from sufficient input views, it is very challenging for them to generalize to new scenes/objects with few samples (Kwon et al.
-2024).
-Recent advancements have explored using neural radiance fields (NeRF) for modeling 3D human avatars (Mildenhall et al.
-2021; Wang et al.
-2022a), typically employing parametric body models to model deformations.
-Some methods (Chen et al.
-2022; Zhao et al.
-2022) use human template models to facilitate generalizable and robust synthesis.
-Though NeRF-based methods have made significant progress in generalizable human rendering, they are less efficient to train and render due to their computationally intensive per-pixel volume rendering process.
-Point-based rendering (Zheng et al.
-2023) has emerged as an efficient alternative to NeRFs, offering significantly faster rendering.
-The recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023) gains popularity for its efficient rasterization-based rendering speed.
-Numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatars (Lei et al.
-2024; Moreau et al.
-2024; Shao et al.
-2024; Hu et al.
-2024b; Kocabas et al.
-2024; Wang et al.
-2024a; Qian et al.
-2024; Hu et al.
-2024a).
+Existing approaches for human avatar generation–both NeRF-based and 3D Gaussian Splatting (3DGS) based–struggle with maintaining 3D consistency and exhibit degraded detail reconstruction, particularly when training with sparse inputs.
+To address this challenge, we propose CHASE, a novel framework that achieves dense-input-level performance using only sparse inputs through two key innovations: cross-pose intrinsic 3D consistency supervision and 3D geometry contrastive learning.
+Building upon prior skeleton-driven approaches that combine rigid deformation with non-rigid cloth dynamics, we first establish baseline avatars with fundamental 3D consistency.
+To enhance 3D consistency under sparse inputs, we introduce a Dynamic Avatar Adjustment (DAA) module, which refines deformed Gaussians by leveraging similar poses from the training set.
+By minimizing the rendering discrepancy between adjusted Gaussians and reference poses, DAA provides additional supervision for avatar reconstruction.
+We further maintain global 3D consistency through a novel geometry-aware contrastive learning strategy.
+While designed for sparse inputs, CHASE surpasses state-of-the-art methods across both full and sparse settings on ZJU-MoCap and H36M datasets, demonstrating that our enhanced 3D consistency leads to superior rendering quality.
+Project page: https://chaseprojectpage.github.io/.
+Figure 1: CHASE.
+We propose an efficient method for creating 3D-consistent animatable avatars from just videos.
+Our method achieve better quality to the most recent SOTA methods [37, 8, 29] in both full and sparse inputs.
+footnotetext: * Equal contributions.footnotetext: †Corresponding Author.footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
+1 Introduction
+Photo-realistic rendering and animation of human bodies is a critical research area with wide-ranging applications in AR/VR, visual effects, virtual try-on, and film production [6].
+Early approaches [23] relied on multi-camera setups to capture high-quality data, requiring extensive computational resources and manual effort.
+While these methods perform well for reconstructing a single scene or object with sufficient input views, they struggle to generalize to new scenes or objects from limited samples [15].
+Recent advancements have explored using neural radiance fields (NeRF) for modeling 3D human avatars [21], typically employing parametric body models to model deformations.
+Some methods [2, 46, 11] use human template models to facilitate generalizable and robust synthesis.
+However, NeRF-based methods are less efficient to train and render due to their computationally intensive per-pixel volume rendering process.
+Point-based rendering [48] has emerged as an efficient alternative to NeRFs, offering significantly faster rendering.
+The recently proposed 3D Gaussian Splatting (3DGS) [12] gains popularity for its fast rendering speed.
+Numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatars [16, 22, 32, 8, 14, 34, 29, 7].
 However, these methods often face challenges in maintaining 3D consistency and producing high-quality reconstructions, particularly with sparse inputs.
-To address the aforementioned issues, we propose CHASE, which is capable of reconstructing 3D Consistent Human Avatars with Sparse inputs via Gaussian Splatting and contrastivE learning.
-We first integrates a skeleton-driven rigid deformation, and a non-rigid cloth dynamics deformation to create human avatar with coarse 3D consistency.
+To address the aforementioned issues, we propose CHASE, which is capable of reconstructing 3D Consistent Human Avatars with Sparse inputs via Gaussian Splatting and contrastivE learning, as shown in Fig.
+1.
+We first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to create a human avatar.
 To enhance 3D consistency under sparse inputs, we utilize the intrinsic 3D consistency of images across different poses within the same person.
 Specifically, for each training pose/image, we select a similar pose/image from the dataset and then adjust the deformed Gaussians using the proposed Dynamic Avatar Adjustment (DAA), an explicit point-based control graph adjustment strategy, to the selected similar pose.
-Then we minimize differences between the rendered image of the adjusted Gaussians and the image corresponding to the selected similar pose, which serves as an additional form of supervision for human avatar.
+Then, we minimize differences between the rendered image of the adjusted Gaussians and the image corresponding to the selected similar pose, which serves as an additional form of supervision for human avatars.
 Additionally, we employ 3D geometry contrastive learning, utilizing features from a 3D feature extractor, to further enhance the global 3D consistency of generated human avatars.
-We conduct extensive experiments on the ZJU-MoCap data (Peng et al.
-2021c), H36M (Ionescu et al.
-2013) and find that CHASE not only achieves better performance in sparse inputs setting but also outperforms other SOTAs in full setting.
-In summary, our work makes the following contributions:
+We conduct extensive experiments on the ZJU-MoCap data [26], H36M [10] and surprisingly find that CHASE outperforms other SOTAs in both full and sparse inputs setting.
+Our work makes the following contributions:
   • We propose an explicit point-based control graph adjustment strategy, which introduces a novel 2D image supervision to 3D human body modeling, enhancing the 3D consistency of human avatars.
-  • We propose 3D geometry contrastive learning to enforce consistency across different representations of the same pose and facilitate effective 3D global understanding.
+  • We propose a 3D geometry contrastive learning to enforce consistency across different representations of the same pose and enhance global 3D understanding.
   • Extensive experiments show that our CHASE achieves SOTA performance quantitatively and qualitatively under full and sparse settings.
-Related Work
-Contrastive Representation Learning
-Contrastive Representation Learning is one of the mainstream self-supervised learning paradigms (Hadsell, Chopra, and LeCun 2006), which learns potential semantics from constructed invariance or equivariance (Rumen et al.
-2022).
-In 3D, PointContrast (Xie et al.
-2020) proposes geometric augmentation to generate positive and negative pairs.
-CrossPoint (Afham et al.
-2022) uses both inter- and intra-modal contrastive learning.
-PointCLIP (Zhang et al.
-2022) achieves image-point alignment by projecting point clouds onto 2D depth images.
-RECON (Qi et al.
-2023) focuses on single- and cross-modal contrastive learning through discriminative contrast (Khosla et al.
-2020) or global feature alignment (Radford et al.
-2021).
-Our CHASE introduces a 3D geometry contrastive learning method across different poses to enforce consistency across different representations of the same pose.
-3D Editing and Deformation
-Traditional deformation methods in computer graphics are typically based on Laplacian coordinates (Gao et al.
-2019; Sorkine and Alexa 2007; Sorkine 2005), Poisson equations (Yu et al.
-2004), and cage-based methods (Yifan et al.
-2020).
+2 Related Work
+2.1 Contrastive Representation Learning
+Contrastive Representation Learning is one of the mainstream self-supervised learning paradigms, which learns potential semantics from constructed invariance or equivariance.
+In 3D, PointContrast [39] proposes geometric augmentation to generate positive and negative pairs.
+CrossPoint [1] uses both inter- and intra-modal contrastive learning.
+PointCLIP [45] achieves image-point alignment by projecting point clouds onto 2D depth images.
+RECON [28] focuses on single- and cross-modal contrastive learning through discriminative contrast [13] or global feature alignment [30].
+Our CHASE introduces a 3D geometry contrastive learning method to enforce consistency across different representations of the same pose.
+2.2 3D Editing and Deformation
+Traditional deformation methods in computer graphics are typically based on Laplacian coordinates [4], Poisson equations [43], and cage-based methods [41].
 However, these methods often rely on implicit and computationally expensive NeRF-based approaches.
-Numerous works (Chen et al.
-2024; Wang et al.
-2024b) have proposed techniques for editing 3D Gaussian Splatting (3DGS)(Kerbl et al.
-2023).
-SuGaR(Guédon and Lepetit 2024) introduces a mesh extraction method that produces meshes from 3DGS, which can then be edited.
-SC-GS (Huang et al.
-2024) proposes deforming Gaussians by transferring the movement of control points.
-Our CHASE employs a novel explicit point-based control graph deformation strategy, which is more intuitive and efficient.
-3D Human Modeling
-Since the high-quality rendering achieved by the seminal work Neural Radiance Fields (NeRF)(Mildenhall et al.
-2021), there has been a surge of research on neural rendering for human avatars(Liu et al.
-2021; Li et al.
-2023; Peng et al.
-2021c).
-Although NeRF is designed for static objects, HumanNeRF (Weng et al.
-2022) extends NeRF to enable capturing dynamic human motion using just a single monocular video.
-Neural Body (Peng et al.
-2021c) associates a latent code with each SMPL (Loper et al.
-2015) vertex to encode appearance, which is then transformed into observation space based on the human pose.
-Furthermore, Neural Actor (Liu et al.
-2021) learns a deformable radiance field with SMPL (Loper et al.
-2015) as guidance and utilizes a texture map to improve the final rendering quality.
-Posevocab (Li et al.
-2023) designs joint-structured pose embeddings to encode dynamic appearances under different key poses, allowing for more effective learning of joint-related appearances.
+Numerous works [3, 47] have proposed techniques for editing 3D Gaussian Splatting (3DGS)[12].
+SuGaR[5] introduces a mesh extraction method that produces meshes from 3DGS, which can then be edited.
+SC-GS [9] proposes deforming Gaussians by transferring the movement of control points.
+Our CHASE employs a novel explicit point-based control graph deformation strategy, which is more efficient than previous methods.
+2.3 3D Human Modeling
+Since the high-quality rendering achieved by the seminal work Neural Radiance Fields (NeRF)[21], there has been a surge of research on neural rendering for human avatars[18, 17, 26, 31].
+Although NeRF is designed for static objects, HumanNeRF [38] extends NeRF to enable capturing dynamic human motion using just a single monocular video.
+Neural Body [26] associates a latent code with each SMPL [19] vertex to encode appearance, which is then transformed into observation space based on the human pose.
+Furthermore, Neural Actor [18] learns a deformable radiance field with SMPL [19] as guidance and utilizes a texture map to improve the final rendering quality.
+Posevocab [17] designs joint-structured pose embeddings to encode dynamic appearances under different key poses, allowing for more effective learning of joint-related appearances.
 However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
-Point-based rendering (Zheng et al.
-2023) has proven to be an efficient alternative to NeRFs for fast inference and training.
-Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS)(Kerbl et al.
-2023) models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending.
-Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatar reconstruction(Lei et al.
-2024; Qian et al.
-2024; Hu et al.
-2024b; Kocabas et al.
-2024).
-Human Gaussian Splatting (Moreau et al.
-2024) showcases 3DGS as an efficient alternative to NeRF.
-SplattingAvatar (Shao et al.
-2024) and GomAvatar (Wen et al.
-2024) extend lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh.
+Point-based rendering [48] has proven to be an efficient alternative to NeRFs for fast inference and training.
+Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS)[12] models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending.
+Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic 3D human avatar reconstruction[16, 29, 8, 14].
+Human Gaussian Splatting [22] showcases 3DGS as an efficient alternative to NeRF.
+SplattingAvatar [32] and GomAvatar [37] extend lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh.
 However, these methods struggle to maintain 3D consistency and produce low-quality reconstructions when applied to human avatar creation with only sparse inputs.
@@
 In contrast, positive pairs include the features of ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the pose adjustments seen in ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{a}}$.
-Preliminaries
-SMPL (Loper et al.
-2015).
+3 Preliminaries
+SMPL [19].
 The SMPL model is a pre-trained parametric human model representing body shape and pose.
@@
 In this work, we apply the Linear Blend Skinning (LBS) algorithm used in SMPL to transform points from a canonical space to a posed space.
-LBS (Sumner, Schmid, and Pauly 2007).
+LBS [33].
 Linear Blend Skinning (LBS) is a weight-based technique that associates each vertex with one or more joints and uses weight values to describe the influence of each joint on the vertex.
@@
 bone transforms) and $\mathcal{X}_{v}\in\mathbb{R}^{N\times 3}$ is the original mesh vertex position.
-3D Gaussian Splatting (3DGS) (Kerbl et al.
-2023).
+3D Gaussian Splatting (3DGS) [12].
 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean.
@@
 $\mathcal{S}$ and $\mathcal{R}$ are stored as the diagonal vector $s\in\mathbb{R}^{N\times 3}$ and a quaternion vector $r\in\mathbb{R}^{N\times 4}$, respectively.
-In rendering novel views, differential splatting as introduced by (Yifan et al.
-2019), involves using a viewing transform $W$ and the Jacobian matrix $J$ of the affine approximation of the projective transformation to compute the transformed covariance matrix: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$.
-The color and opacity at each pixel are computed from the Gaussian’s representations: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ The blending of $N$ ordered points overlapping a pixel is given by the formula:
-$\mathcal{C}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ (2)
-where $c_{i}$, $\alpha_{i}$ represent the density and color of this point computed by a 3D Gaussian $G$ with covariance $\Sigma$ multiplied by an optimizable per-point opacity and SH color coefficients.
-Method
+In rendering novel views, differential splatting as introduced by [42], involves using a viewing transform $W$ and the Jacobian matrix $J$ of the affine approximation of the projective transformation to compute the transformed covariance matrix: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$.
+The color and opacity at each pixel are computed from the Gaussian’s representations: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ The blending of $N$ ordered points overlapping a pixel is given by the formula: $\mathcal{C}=\sum_{i\in N}c_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{i}),$ where $c_{i}$, $\alpha_{i}$ represent the density and color of this point computed by a 3D Gaussian $G$ with covariance $\Sigma$ multiplied by an optimizable per-point opacity and SH color coefficients.
+4 Method
 We illustrate the pipeline of our CHASE in Fig.
-1.
+2.
 The inputs include images $X=\{x_{i}\}^{N}_{i=1}$ obtained from monocular videos, fitted SMPL parameters $P=\{p_{i}\}^{N}_{i=1}$, and foreground masks $M=\{m_{i}\}^{N}_{i=1}$ of images.
 CHASE optimizes 3D Gaussians in canonical space, which are then be deformed to match the observation space and be rendered with a given camera view.
-Non-rigid and Rigid Deformation
-Inspired by (Weng et al.
-2022; Qian et al.
-2024), we deform 3D Gaussians from canonical space ${\mathcal{G}_{c}}$ to observation space ${\mathcal{G}_{o}}$ by integrating a rigid articulation with a non-rigid transformation.
-We employ a non-rigid deformation network that takes the canonical positions $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}c}$ and a pose latent code which encodes SMPL pose $p_{i}$ using a lightweight hierarchical pose encoder (Mihajlovic et al.
-2021) as input.
+4.1 Non-rigid and Rigid Deformation
+Inspired by [38, 29], we deform 3D Gaussians from canonical space ${\mathcal{G}_{c}}$ to observation space ${\mathcal{G}_{o}}$ by integrating a rigid articulation with a non-rigid transformation.
+We employ a non-rigid deformation network that takes the canonical positions $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}c}$ and a pose latent code which encodes SMPL pose $p_{i}$ using a lightweight hierarchical pose encoder [20].
 The network then outputs the offsets for various parameters of the 3D Gaussians $\mathcal{G}_{c}$: $\Delta{(\mathcal{X},\mathcal{C},\alpha,s,r)}$.
@@
 where $\mathcal{R}$ is the matrix representations of rotation.
-Figure 2: For each training pose/image, we select similar poses/images from the dataset and then adjust the deformed Gaussians using Dynamic Avatar Adjustment (DAA).
-By minimizing the discrepancy between the rendered image of the adjusted avatar and the selected similar pose image, we introduce an additional supervision, thereby enhancing the creation of animatable avatars.
-Dynamic Avatar Adjustment
+Figure 3: For each training pose/image, we select similar poses/images from the dataset and then adjust the deformed Gaussians using DAA.
+By minimizing the difference between the rendered image of the adjusted avatar and the selected similar pose image, we introduce additional supervision, thereby refining the creation of photo-realistic and animatable avatars.
+4.2 Dynamic Avatar Adjustment
 To address extremely sparse inputs, we leverage the intrinsic 3D consistency of human avatars across different poses/images, as shown in Fig.
-2.
-Specifically, for each training pose/image, we select a similar pose $p^{a}_{i}$ with its paired image $x_{i}^{a}$ from the dataset and then use a dense motion field $F_{adj}$ as an additional adjustment to transform deformed Gaussians ${\mathcal{G}_{o}}$ into adjusted Gaussians ${\mathcal{G}_{a}}$, aligning them with the selected pose/image ($p_{i}^{a}$/$x_{i}^{a}$).
+3.
+Though the same pose may exhibit slight variations in non-rigid deformation, these differences occupy only a small number of pixels from a whole-body perspective and thus have minimal impact on the overall body reconstruction, which we further demonstrate in the experimental Section.
+5.3.
+Specifically, for each training pose/image, we select a similar pose $p^{a}_{i}$ with its paired image $x_{i}^{a}$ by computing the orientation and limb angle difference provided by SMPL [19] model from the dataset.
+Then we use a dense motion field $F_{adj}$ as an additional adjustment to transform deformed Gaussians ${\mathcal{G}_{o}}$ into adjusted Gaussians ${\mathcal{G}_{a}}$, aligning them with the selected pose/image ($p_{i}^{a}$/$x_{i}^{a}$).
 In this way, we successfully introduce an additional 2D image supervision, improving the 3D consistency of human avatars.
-To achieve precise control of the 3D Gaussians, we sample 6,890 points from the SMPL model (Loper et al.
-2015) as our sparse control points in canonical space.
+To achieve precise control of the 3D Gaussians, we sample 6,890 points from the SMPL model [19] as our sparse control points in canonical space.
 Then, we obtain the dense motion field using LBS by locally inheriting the LBS weights from neighboring control points.
@@
 $\mathcal{R}_{a}$ $=F_{adj1:3,1:3}\mathcal{R}_{o}.$ (11)
-3D Geometry Contrastive learning
-Inspired by the success of contrastive learning in image processing and static point cloud analysis, we advocate for adopting 3D geometry contrastive learning to ensure 3D consistency during animation.
-To this end, we treat the 3D Gaussians as a 3D point cloud and use DGCNN (Wang et al.
-2019) as the feature extractor.
+4.3 3D Geometry Contrastive learning
+Inspired by the success of contrastive learning in 2D image processing and static point cloud analysis, we advocate for adopting 3D geometry contrastive learning to ensure 3D consistency of avatars.
+We treat the 3D Gaussians as a 3D point cloud and use DGCNN [36] as the feature extractor.
+DGCNN is typically trained on general point cloud datasets, which helps it learn geometric structures.
 The point cloud feature extractor processes the positions of the 3D Gaussians in the observation space ${\mathcal{G}_{o}}$, the adjusted Gaussians ${\mathcal{G}_{a}}$, and ${\mathcal{G}_{o}^{\prime}}$, which is deformed from the canonical space to match the selected pose $p_{i}^{a}$, and outputs their features, creating intermediate graph features to capture global geometric information better.
@@
 In the invariant space, we aim to maximize the similarity between $f_{a}$ and $f_{o}^{\prime}$, denoted as $\text{D}_{positive}$, and minimize the similarity between $f_{a}$ and $f_{o}$, denoted as $\text{D}_{negative}$.
+This feature-level contrastive learning encourages the model to capture pose-aware knowledge, enabling it to differentiate between subtle pose variations while maintaining geometric consistency.
 Therefore, we compute the 3D geometry contrastive loss $\mathcal{L}_{contrastive}$ as:
@@
 $\mathcal{L}_{contrastive}=\max(0$ $,\text{D}_{positive}-\text{D}_{negative}).$ (14)
-Optimization
+4.4 Optimization
+We begin by randomly sampling 50k points from the surface of the SMPL mesh to initialize 3D Gaussians in the canonical space.
 Color MLP.
-Following (Qian et al.
-2024), we use the inverse rigid transformation to canonicalize the viewing direction: $\hat{d}=T_{1:3,1:3}^{-1}d$, where $T$ and $d$ is the forward transformation matrix defined in LBS and viewing direction, respectively.
+Following [29], we use the inverse rigid transformation to canonicalize the viewing direction: $\hat{d}=T_{1:3,1:3}^{-1}d$, where $T$ and $d$ is the forward transformation matrix defined in LBS and viewing direction, respectively.
 Theoretically, canonicalizing viewing direction also promotes consistency of the specular component of canonical 3D Gaussians under rigid transformations.
 Pose correction.
-Following (Qian et al.
-2024), SMPL (Loper et al.
-2015) parameter fittings from images can be inaccurate.
+Following [29], SMPL [19] parameter fittings from images can be inaccurate.
 We additionally optimize the per-sequence shape parameter and per-frame translation, global rotation, and local joint rotations.
@@
 We compute these losses on both images rendered from the deformed Gaussians $\mathcal{G}_{o}$ and the adjusted Gaussians $\mathcal{G}_{a}$ with their corresponding ground truth images.
-Additionally, we include a skinning weight regularization loss $\mathcal{L}_{skin}$, as well as isometric regularization losses for both position and covariance, $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$, respectively, following (Qian et al.
-2024).
+Additionally, we include a skinning weight regularization loss $\mathcal{L}_{skin}$, as well as isometric regularization losses for both position and covariance, $\mathcal{L}_{isopos}$ and $\mathcal{L}_{isocov}$, following [29].
 We also incorporate a 3D geometry contrastive loss $\mathcal{L}_{contrastive}$:
@@
 where $\lambda$’s are loss weights. For further details of the loss definition and respective weights, please refer to the Supp.Mat.
-Figure 3: Qualitative Comparison on ZJU-MoCap (Peng et al.
-2021c).
+Figure 4: Qualitative Comparison on ZJU-MoCap [26].
 We present results for full and sparse inputs (5% of the full inputs) on the ZJU-MoCap dataset.
 Results show that our CHASE can produce realistic details with both full and sparse inputs, while other approaches struggle to generate smooth details.
-Table 1: Quantitative Results on ZJU-MoCap (Peng et al.
-2021c).
+Figure 5: Qualitative Comparison on H36M [10] with sparse inputs.
+We demonstrate that our method effectively produces realistic details for novel pose in both rendered images and geometry, whereas other approach struggles to achieve smooth details.
+Table 1: Quantitative Results on ZJU-MoCap [26].
 CHASE achieves state-of-the-art performance across every method.
@@
 […]
-NeuralBody (Peng et al. 2021c) 29.07 0.962 52.29 1.5
-Ani-NeRF (Peng et al. 2021a) 29.17 0.961 51.98 1.1
-HumanNeRF (Weng et al. 2022) 30.24 0.968 31.73 0.3
-MonoHuman (Yu et al. 2023) 29.38 0.964 37.51 0.1
-3DGS-Avatar (Qian et al. 2024) 30.62 0.965 30.28 50
-GauHuman (Hu et al. 2024b) 30.79 0.960 32.73 180
-GoMAvatar (Wen et al. 2024) 30.37 0.969 32.53 43
-Ours 30.81 0.970 27.48 50
-3DGS-Avatar* (Qian et al. 2024) 29.98 0.957 40.01 50
-GauHuman* (Hu et al. 2024b) 30.35 0.957 35.68 180
-GoMAvatar* (Wen et al. 2024) 30.01 0.958 42.88 43
-Ours* 30.48 0.969 29.94 50
-Table 2: Quantitative Results on H36M (Ionescu et al. 2013). Our CHASE outperforms current SOTA methods in both full and sparse settings
+NeuralBody [26] 52.29 29.07 0.962 1.5
+Ani-NeRF [25] 51.98 29.17 0.961 1.1
+HumanNeRF [38] 31.73 30.24 0.968 0.3
+MonoHuman [44] 37.51 29.38 0.964 0.1
+DVA [31] 37.74 29.45 0.956 17
+InstantAvatar [11] 64.41 29.73 0.938 4.2
+3DGS-Avatar [29] 30.28 30.62 0.965 50
+GauHuman [8] 32.73 30.79 0.960 180
+GoMAvatar [37] 32.53 30.37 0.969 43
+Ours 27.48 30.81 0.970 50
+3DGS-Avatar* [29] 40.01 29.98 0.957 50
+GauHuman* [8] 35.68 30.35 0.957 180
+GoMAvatar* [37] 42.88 30.01 0.958 43
+Ours* 29.94 30.48 0.969 50
+5 Experiment
+5.1 Dataset
+ZJU-MoCap [26].
+This dataset features multi-view videos captured by 21 cameras, with human poses recorded using a marker-less motion capture system.
+For our experiments, we selected six sequences (377, 386, 387, 392, 393, 394).
+Following the protocol established by HumanNeRF [38] and 3DGS-Avatar [29], we use a single camera for training and the remaining cameras for evaluation.
+The foreground masks, camera, and SMPL parameters provided by the data set are used for evaluation purposes.
+We simulate sparse inputs by sampling every 20th frame from the video, between start_frame and end_frame, which corresponds to using 5% of the total frames.
+We also specify in the Supp.Mat which images are selected for training.
+H36M [10].
+H36M is another widely used dataset for human avatar research, comprising multi-view videos from four cameras and human poses captured via a marker-based motion capture system.
+We conducted experiments on sequences from subjects S1, S5, S6, S7, S8, S9, and S11, selecting representative actions and dividing the videos into training and test frames.
+Adhering to the protocol set by ARAH [35], we use three cameras, [54138969, 55011271, 58860488], for training and the remaining camera, [60457274], for testing, and follow their preprocessing steps.
+We use the SMPL parameters and foreground humans following [25].
+We apply the same approach as mentioned above to simulate sparse inputs in this dataset.
+Table 2: Quantitative Results on H36M [10]. Our CHASE outperforms current SOTA methods in both full and sparse settings
 Training Poses Novel Poses
 […]
-NARF (Noguchi et al. 2021) 23.00 0.898 22.27 0.881
-NeuralBody (Peng et al. 2021c) 22.89 0.896 23.09 0.891
-Ani-NeRF (Peng et al. 2021a) 23.00 0.890 22.55 0.880
-ARAH (Wang et al. 2022b) 24.79 0.918 23.42 0.896
-3DGS-Avatar (Qian et al. 2024) 32.89 0.982 32.50 0.983
+NARF [24] 23.00 0.898 22.27 0.881
+NeuralBody [26] 22.89 0.896 23.09 0.891
+Ani-NeRF [25] 23.00 0.890 22.55 0.880
+ARAH [35] 24.79 0.918 23.42 0.896
+3DGS-Avatar [29] 32.89 0.982 32.50 0.983
 Ours 33.29 0.984 32.93 0.982
-3DGS-Avatar* (Qian et al. 2024) 32.48 0.976 32.17 0.981
+3DGS-Avatar* [29] 32.48 0.976 32.17 0.981
 Ours* 32.91 0.983 32.64 0.982
-Experiment
-Dataset
-ZJU-MoCap (Peng et al.
-2021c).
-This dataset features multi-view videos captured by 21 cameras, with human poses recorded using a marker-less motion capture system.
-For our experiments, we selected six sequences (377, 386, 387, 392, 393, 394).
-Following the protocol established by HumanNeRF (Weng et al.
-2022) and 3DGS-Avatar (Qian et al.
-2024), we use a single camera for training and the remaining cameras for evaluation.
-The foreground masks, camera, and SMPL parameters provided by the data set are used for evaluation purposes.
-We use 5% of the dataset to simulate sparse inputs.
-H36M (Ionescu et al.
-2013).
-H36M is another widely used dataset for human avatar research, comprising multi-view videos from four cameras and human poses captured via a marker-based motion capture system.
-It features multiple subjects performing a variety of complex actions.
-We conducted experiments on sequences from subjects S1, S5, S6, S7, S8, S9, and S11, selecting representative actions and dividing the videos into training and test frames.
-Adhering to the protocol set by ARAH (Wang et al.
-2022b), we use three cameras, [54138969, 55011271, 58860488],for training and the remaining camera, [60457274],for testing, and follow their preprocessing steps.
-We use the SMPL parameters and foreground humans following (Peng et al.
-2021a).
-Comparison with State-of-the-art Methods
-We compare our CHASE with various SOTA methods for human avatars, including NerF-based methods such as NeuralBody (Peng et al.
-2021c), Ani-NeRF (Peng et al.
-2021a), HumanNeRF (Weng et al.
-2022), and MonoHuman (Yu et al.
-2023), and 3DGS-based methods such as 3DGS-Avatar (Qian et al.
-2024), GauHuman (Hu et al.
-2024b), and GoMAvatar (Wen et al.
-2024) under monocular setup on ZJU-MoCap (Peng et al.
-2021c).
+5.2 Comparison with State-of-the-art Methods
+We compare our CHASE with various SOTA methods for human avatars, including NerF-based methods such as NeuralBody [26], Ani-NeRF [25], HumanNeRF [38], and MonoHuman [44], and 3DGS-based methods such as 3DGS-Avatar [29], GauHuman [8], and GoMAvatar [37] under monocular setup on ZJU-MoCap [26].
 The quantitative results are shown in Tab.
@@
 Overall, our proposed CHASE achieves the best performance in terms of PSNR, SSIM, and LPIPS with both full and sparse inputs.
-Our CHASE shows only a small performance drop when using only 5% of the data and significantly outperforms SOTA methods on LPIPS.
+Notably, our CHASE shows only a small performance drop when using only 5% of the data, especially in LPIPS which is more informative than the other two metrics in our setting [29].
+In fact, it declines less compared to other methods and even surpasses their performance with 100% of the data.
 It is evidence that our method successfully maintains 3D consistency even with sparse inputs.
 Qualitative comparisons on novel view synthesis are shown in Fig.
-3.
+4.
 We observe that our method preserves more details compared to other SOTA methods.
 They often struggle to maintain 3D consistency and deliver suboptimal detail reconstruction in human avatar modeling, particularly when only sparse inputs are available.
-For H36M (Ionescu et al.
-2013), we report the quantitative results against NerF-based methods such as NARF (Noguchi et al.
-2021), NeuralBody (Peng et al.
-2021c), Ani-NeRF (Peng et al.
-2021a) and ARAH (Wang et al.
-2022b), and 3DGS-based methods such as 3DGS-Avatar (Qian et al.
-2024) in Tab.
+Please see our project website and supplementary material for more visualization.
+For H36M [10], we report the quantitative results against NeRF-based methods such as NARF [24], NeuralBody [26], Ani-NeRF [25] and ARAH [35], and 3DGS-based methods such as 3DGS-Avatar [29] in Tab.
 2.
 Our CHASE significantly outperforms these methods with sparse inputs, showing that our CHASE generalizes well to novel poses with sparse inputs and reconstructs human avatars with better appearance and geometry detail.
-For qualitative comparisons on novel pose synthesis, please refer to the Supp.Mat.
-Table 3: Ablation Study on ZJU-MoCap (Peng et al. 2021c). We both show the result from full input (top group) and sparse input (bottom group).
+For qualitative comparisons on novel pose synthesis, as shown in Fig.
+5, our method generalizes well to novel pose with just sparse inputs (only 5% of origin data) and reconstruct human avatars with better appearance and geometry detail.
+Table 3: Ablation Study on ZJU-MoCap [26]. We both show the result from full input (top group) and sparse input (bottom group).
 […]
@@
 w/o DAA 30.78 0.970 27.83 50
+Top-3 30.80 0.970 27.73 50
+Top-5 30.79 0.970 27.74 50
+Top-10 30.76 0.970 27.82 50
 Full model 30.81 0.970 27.48 50
@@
 w/o DAA 30.42 0.968 30.17 50
-Full model 30.41 0.969 29.94 50
-Figure 4: Ablation Study on Dynamic Avatar Adjustment, which enhances multi-view consistency, hence improving rendering quality.
+Top-3 30.44 0.969 29.98 50
+Top-5 30.43 0.968 30.19 50
+Full model 30.48 0.969 29.94 50
+Figure 6: Ablation Study on DAA, which enhances multi-view 3D consistency, hence improving the overall rendering quality.
 Figure 7: Ablation Study on 3D geometry contrastive learning, which removes the artifacts on highly articulated poses..
-Ablation Study
-In this section, we conduct ablation experiments using the ZJU-MoCap (Peng et al.
-2021c) dataset with both full inputs and sparse inputs to evaluate the effectiveness of our proposed modules.
+5.3 Ablation Study
+In this section, we conduct ablation experiments using the ZJU-MoCap [26] dataset with both full inputs and sparse inputs to evaluate the effectiveness of our proposed modules.
 We also conduct experiments to evaluate different backbones in 3D geometry contrastive learning.
@@
 Non-rigid deformation.
-Non-rigid Deformation is designed for complex cloth deformation.
+Non-rigid Deformation is designed for complex deformations, such as cloth bending and stretching.
 As shown in Tab.
@@
 Dynamic Avatar Adjustment.
-As shown in Tab.3, incorporating Dynamic Avatar Adjustment (DAA) results in our full model outperforming the baseline in terms of LPIPS.
-We argue that LPIPS is particularly informative compared to other metrics because it addresses the challenge of accurately reproducing ground-truth appearance from novel views in a monocular setting(Qian et al.
-2024; Yang et al.
-2024).
+As shown in Tab.3, incorporating DAA results in our full model outperforming the baseline in terms of LPIPS which is particularly informative compared to other metrics in our setting [29, 40].
 Fig.
-4 illustrates that Dynamic Avatar Adjustment serves as an effective 2D image supervision technique for 3D human body modeling, enhancing 3D consistency and reducing artifacts while improving multi-view consistency.
+6 illustrates that DAA serves as an effective 2D image supervision for 3D human body modeling, enhancing 3D consistency and reducing artifacts while improving multi-view consistency.
+Additionally, DAA reduces artifacts and inaccuracies in the geometry caused by sparse inputs, further enhancing the robustness and visual realism of the model in practical applications.
+Similar (pose/image) selection.
+Our method selects a similar (pose/image) to supervise the generated avatar as described in Section 4.2.
+Ablation studies in Tab.
+3 show that even when selecting less similar (poses/images), our approach significantly improves performance.
+Top-n refers to choosing the n-th most similar pose/image.
+These results highlight CHASE ’s effectiveness in handling limited data or irregular large-scale human movements.
 3D geometry contrastive learning.
@@
 3 and Fig.
-5.
-We can find the full model (w/ contrastive) preserves finer details and provides a more realistic and detailed reconstruction of clothing and other complex surfaces, demonstrating that 3D Geometry Contrastive Learning enhances 3D consistency.
+7.
+We can find the full model (w/ contrastive) preserves finer details and provides a more realistic and detailed reconstruction of clothing, demonstrating that 3D geometry contrastive learning enhances 3D consistency.
 Backbone for 3D contrastive learning.
 In Tab.
-3, We show the ablation study on different backbones, including PointNet (Qi et al.
-2017) (pointnet) and DGCNN (Wang et al.
-2019) (full model).
+3, we show the ablation study on different backbones, including PointNet [27] (pointnet) and DGCNN [36] (full model).
 This indicates that DGCNN captures local geometric details better by building a dynamic graph structure, while PointNet relies on global feature learning, which may cause some local information to be missing.
-Conclusion
+6 Conclusion
 In this paper, we present CHASE, a 3D-consistent human modeling framework utilizing Gaussian Splatting with both full and sparse inputs.
@@
 2).
-Our CHASE need to find similar pose from dataset.
-When it comes to monocular videos featuring large-scale movements and diverse actions, it is hard to find pairs that qualify the requirements, which may impact performance.
+CHASE performs less effectively in reconstructing humans with complex clothing, as similar poses may still exhibit substantial differences in non-rigid deformations.
+This remains a challenge we plan to address in future work.